# Evaluating and improving wave and non-wave stress parametrisations for oceanic flows

### v1

@@ 4.3 Steady flow in two dimensions
For two-dimensional tests with a steady flow we compare the computed stress with the parametrisations $F_{\text{Bell2d}}$ and $F_{\text{KLP}}$ from Section 2. In particular, tests with low Froude numbers $Fr\lesssim 1$ are compared with $F_{\text{Bell2d}}$ and tests with high Froude numbers $Fr\gg 1$ are compared with $F_{\text{KLP}}$.
Figure 5a shows our comparison between the simulated data and $F_{\text{Bell2d}}$ for Froude numbers in the range $0.1-1$. We see that although the scaling used in Bell’s theory is generally accurate, $F_{\text{Bell2d}}$ typically underestimates the stress by about $50\%$, and even more so as the Froude number approaches 1. Hence, Bell’s parametrisation should ideally be adjusted to depend on $Fr$. In this regard, one could define a scaled parametrisation
@@
for constants $A_{1}$, $B_{1}$, and $Fr\lesssim 1$. Note that the linear trend line in Figure 5a corresponds to the values $A_{1}=0.71$ and $B_{1}=0.46$. However, a parametrisation of the form (31) still does not describe the full behaviour of the stress in the regime $Fr\lesssim 1$. Namely, in Figure 5b we also see a non-linear relationship between the stress and latitude. This reinforces that, even for low Froude numbers, non-wave and non-linear effects manifest that are not accounted for by Bell’s theory. We also remark that the trend line in Figure 5a is the opposite to what one gets with energy flux calculations. That is, as the Froude number increases, one reaches a saturation of lee waves and Bell’s parametrisation (for energy flux) becomes an overestimate (see (17)).
Figures 5c–5e then show our comparison between the simulated data and $F_{\text{KLP}}$ for large Froude numbers and $f=0$. Here, the approximately linear scaling of $N$ and $U$ in $F_{\text{KLP}}$ seems to be very accurate (Figures 5c and 5d). On the other hand, the height scaling of $h_{0}^{2}H/(H-h_{0})$ overestimates the stress as $h_{0}$ gets large (Figure 5e). Our results indicate that a simpler scaling than that in $F_{\text{KLP}}$ is more accurate. In particular, if we set
$F_{\text{2d}}=C_{1}Nh_{0}^{2}U$ (32)
@@
We also ran a few tests with rotation added to the high Froude number simulations. In this case, the stress was unstable and increased by several fold. Upon analysing the simulation data, we found that for these cases significant mixing occurred downslope, and gradually reduced the downstream stratification to zero. This phenomena is reminiscent of atmospheric downslope windstorms, which can occur near large-scale topography on land (see (e.g., Peltier and Clark, 1979; Scinocca and McFarlane, 2000). However, this dramatic effect was only observed in our two-dimensional tests, since in three-dimensions the presence of a lateral flow helps to stabilise the stratification. Therefore, such effects are mostly unphysical and we have not attempted to parametrise them here.
Figure 5: A comparison of the stress predicted by the Bell (1975) and KLP2010 parametrisations with our simulation data for two-dimensional steady flow. The default parameters are as in Table 1 and error bars are shown whenever the uncertainty in the stress was greater than 10%.
@@ 4.4 Steady flow in three dimensions
For steady flow tests in three-dimensions, we again compare simulations with low Froude numbers $Fr\lesssim 1$ to the Bell (1975) theory. However, for tests with high Froude numbers $Fr\gg 1$, it does not make sense to compare our results directly to the KLP2010 theory as this theory was only derived in the two-dimensional case. Thus, for high Froude numbers, we instead focus on how the stress scales with different parameters.
Figure 6 shows a comparison between the stress from our simulations and $F_{\text{Bell3d}}$ when the parameters $h_{0}$, $N$, $U$, $f$ and $W$ are varied. For tests varying $h_{0}$ and $N$ (Figure 6a) we find, similar to the two-dimensional case, that $F_{\text{Bell3d}}$ underestimates the stress but still gives a mostly accurate scaling (i.e. $F_{\text{3d}}\propto Nh_{0}^{2}$). However, for tests that vary $U$, $f$ and $W$ (Figure 6b), we see that $F_{\text{Bell3d}}$ significantly underestimates the stress as the Rossby number
$Ro\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{U}{|f|W}$ (33)
@@
where $\tilde{K}=W\sqrt{k^{2}+\ell^{2}}$ and $(kW,\ell W)=(\tilde{K}\cos\theta,\tilde{K}\sin\theta)$. From (4.4) we see that as $Ro$ increases so too does the lee wave stress. Therefore, a possible explanation for the discrepancy between $F_{\text{Bell3d}}$ and the simulation data is that as $Ro$ decreases, fewer lee waves are formed and the stress is instead generated by other non-linear motions such as partial blocking (see e.g., Trossman et al., 2016; Klymak et al., 2021). To parametrise this phenomenon, we note that from Figure 6b the ratio $F_{\text{Bell3d}}/F_{3d}$ has a clear linear dependence on the ratio $fW/U$. In particular, our data points give a correlation coefficient between $F_{\text{Bell3d}}/F_{3d}$ and $fW/U$ of $R^{2}=0.89$. This suggests that we should use the scaling
@@
Figure 7 shows the results of our tests for the high Froude number $Fr\gg 1$ regime. In this case, we find that the stress deviates significantly from the approximate $Nh_{0}^{2}WU$ scaling found in other regimes. Instead we find that the stress obeys a quadratic law $F_{3d}\propto h_{0}WU^{2}$, similar to what one expects in the atmosphere (e.g., Lott and Miller, 1997). In fact, the simple parametrisation $F_{3d}=h_{0}WU^{2}$ works remarkably well for our choice of parameters. The stress then appears to be mostly independent of $N$ and $f$ (Figures 7d and 7e); only increasing or decreasing by small amounts as these parameters are varied. We note however, that for larger values of $f$, our computed value of the stress has a large uncertainty. These results can be directly compared to recent work of Klymak et al. (2021) which involved a similar analysis but with many, randomly distributed obstacles. Klymak et al. also observed this $f$ independence for the stress but an approximately linear dependence on $N$ for high Froude numbers. Therefore, we do not expect the stress to remain independent of $N$ for more complicated non-isolated topography.
Figure 7: Topographic stress for three-dimensional steady flows with high Froude number $(Fr\gg 1)$. In panels (a)–(c), the drag from the simulation data is compared with both a linear $Nh_{0}^{2}WU$ and quadratic $h_{0}WU^{2}$ scaling. The default parameters are as in Table 1. Note that in panels (a)–(d) the uncertainty in the computed stress was always less than 10% so the (small) error bars have been omitted.
@@ 4.5 Mixed flow in two dimensions
We now analyse our results for mixed flows with a horizontal velocity of the form (20) and stress of the form (27). In what follows we let $\widetilde{F}_{m}$ and $\widetilde{F}_{p}$ be the values for $F_{m}$ and $F_{p}$ if $U_{\text{tidal}}$ or $U_{\text{mean}}$ are set to be zero respectively. This lets us discern the impact of a tidal flow on a steady flow and vice versa. Note that values for $\widetilde{F}_{m}$ and $\widetilde{F}_{p}$ are obtained using simulations as in the purely steady or tidal flow cases. In Figure 8 we show our computed values of $F_{m}/\widetilde{F}_{m}$ and $F_{p}/\widetilde{F}_{p}$ for a range of parameters. When the Froude number is small ($Fr\lesssim 1$) these ratios are also compared to the expected ratio from Bell’s parametrisations defined in Sections 22.2 and 22.3 (Figures 8e and 8f).
@@
For low Froude numbers (Figures 8e and 8f) we also observe a dependence of $F_{p}/\widetilde{F}_{p}$ on the parameters $U_{\text{mean}}$ and $W$. In particular, $F_{p}/\widetilde{F}_{p}$ notably exceeds 1 for large values of $U_{\text{mean}}$ (Figure 8e), and small values of $W$ (Figure 8f). This is likely linked to the fact that the classical lee wave frequency $kU_{\text{mean}}$ depends explicitly on $U_{\text{mean}}$ and implicitly on $W$. In particular, with the Fourier convention in (5), the wavenumber spectrum for a Gaussian hill $h(x)=h_{0}e^{-x^{2}/(2W^{2})}$ is
@@
It appears difficult however, to convert these observations into an accurate analytic parametrisation for $F_{p}$.
@@ 4.6 Mixed flow in three dimensions
Finally we consider mixed flows in three dimensions, with our results for this regime shown in Figure 9. Note that we use the same notation as in Section 34.5).
Like the two-dimensional case, we also find that $F_{m}/\widetilde{F}_{m}\approx 1$ and $F_{p}/\widetilde{F}_{p}\approx 1$ for the majority of tests. In fact, this approximation holds true even more often than the two-dimensional case (cf. Figure 8). Of particular note is the far less dramatic increase in $F_{p}/\widetilde{F}_{p}$ near the critical latitude $|f|\approx\omega$. There is however, a distinct increase in the ratio $F_{m}/\widetilde{F}_{m}$ as $h_{0}$ increases, which was not observed in the two-dimensional setting.
@@
$F_{3d}=F_{m}+F_{p}\cos(\omega t+\phi)$ (38)
with $\phi\approx 0$. By comparison, in two dimensions we found that $\phi$ was much larger as $|f|$ increased (Figure 8g). Such an effect is also in contrast to the purely tidal flow case, whereby the phase $\phi$ approaches $-\pi/2$ for large $|f|$.
@@ 5 Discussion and summary of suggested parametrisations
In the previous section, we tested existing topographic stress parametrisations for oceanic flows against a suite of simulations. Particular parametrisations of interest were those of Shakespeare et al. (2020) for tidal flows, as well as Bell (1975) and Klymak et al. (2010) for steady flows. Spanning a wide parameter space for an isolated Gaussian hill, we found that in many cases these parametrisations proved accurate, sometimes even when tested beyond their assumptions. In other cases, it was clear that an alternate parametrisation is required.
We also studied the case of a mixed flow in Sections 34.5 and 34.6. Here, we primarily compared our simulation data against the naive parametrisation which assumes that the tidal and steady components of a flow are independent. Our data yielded the cases for which this naive parametrisation was accurate, and for which cases a more sophisticated parametrisation needs to be developed.
@@ 5.1 Suggested tidal flow parametrisations
For a purely tidal flow, we found that the parametrisations $F_{\text{SAH2d}}$ (Equation (2.1.2)) and $F_{\text{SAH3d}}$ (Equation (2.1.2) due to Shakespeare et al. (2020, 2021) were very accurate, mainly just requiring a small modification for large hill heights. In particular, for a purely oscillatory (tidal) flow in two dimensions, we suggest setting
@@
noting that this parametrisation may overestimate the stress when $h_{0}\approx H$.
@@
for large hill widths $W$, where $F_{\text{SAH3d}}$ is as in (2.1.2). From our testing, $W=10000\ \mathrm{m}$ was a sufficiently large width to notice this effect (with $h_{0}=100\ \mathrm{m}$). For more modest values of $W$, we instead suggest setting $F_{3d}=F_{\text{SAH3d}}$.

### v3

@@ 4.3 Steady flow in two dimensions
Figure 5: A comparison of the stress predicted by the Bell (1975) and KLP2010 parametrisations with our simulation data for two-dimensional steady flow. Here, NN denotes the buoyancy frequency, $h_{0}$ the hill height, $U$ the flow speed, and $f$ the Coriolis parameter. Note that for Figures (a) and (b), the ratio between the stress as predicted the Bell (1975) theory, $F_{\text{Bell2d}}$, and the stress output from the simulations, $F_{\text{2d}}$, is plotted. On the other hand, in Figures (c) to (e) the numerical values of the stress predicted by KLP2010 parametrisation, $F_{\text{KLP}}$, is shown. The default parameters are as in Table 1 and error bars are shown whenever the uncertainty in the stress was greater than 10%.
For two-dimensional tests with a steady flow we compare the computed stress (denoted $F_{\text{2d}}$) with the parametrisations $F_{\text{Bell2d}}$ and $F_{\text{KLP}}$ from Section 2. In particular, tests with low Froude numbers $Fr\lesssim 1$ are compared with $F_{\text{Bell2d}}$ and tests with high Froude numbers $Fr\gg 1$ are compared with $F_{\text{KLP}}$.
Figure 5a shows our comparison between the simulated data and $F_{\text{Bell2d}}$ for Froude numbers in the range 0.110.1-1. We see that although the scaling used in Bell’s theory is generally accurate, $F_{\text{Bell2d}}$ typically underestimates the stress by about $50\%$, and even more so as the Froude number approaches 1. Hence, Bell’s parametrisation should ideally be adjusted to depend on $Fr$. In this regard, one could define a scaled parametrisation
@@
for constants $A_{1}$, $B_{1}$, and $Fr\lesssim 1$. Note that the linear trend line in Figure 5a corresponds to the values $A_{1}=0.71$ and $B_{1}=0.46$. However, a parametrisation of the form (34) still does not describe the full behaviour of the stress in the regime $Fr\lesssim 1$. Namely, in Figure 5b we also see a non-linear relationship between the stress and latitude. This reinforces that, even for low Froude numbers, non-wave and non-linear effects manifest that are not accounted for by Bell’s theory. We also remark that the trend line in Figure 5a is the opposite to what one gets with energy flux calculations. That is, as the Froude number increases, one reaches a saturation of lee waves and Bell’s formula for energy flux becomes an overestimate (see (17)).
Figures 5c–5e then show our comparison between the simulated data and $F_{\text{KLP}}$ for large Froude numbers and $f=0$. Here, the approximately linear scaling of buoyancy frequency $N$ and flow speed $U$ in $F_{\text{KLP}}$ seems to be very accurate (Figures 5c and 5d). On the other hand, the height scaling of $h_{0}^{2}H/(H-h_{0})$ overestimates the stress as the hill height $h_{0}$ gets large (Figure 5e). Our results indicate that a simpler scaling than that in $F_{\text{KLP}}$ is more accurate. In particular, if we parametrise the stress as
$\text{Stress}=C_{1}Nh_{0}^{2}U,$ (35)
@@
We also ran a few tests with rotation added to the high Froude number simulations. In this case, the stress was unstable and increased by several fold. Upon analysing the simulation data, we found that for these cases significant mixing occurred downslope, and gradually reduced the downstream stratification to zero. This phenomena is reminiscent of atmospheric downslope windstorms, which can occur near large-scale topography on land (see e.g., Peltier and Clark, 1979; Scinocca and McFarlane, 2000). However, this effect was far less dramatic in our three-dimensional simulations, where the presence of a lateral flow helps to stabilise the stratification. Such effects are thus likely to be rare in practice, mostly restricted to ridge-like structures. Rather than attempting to parametrise this non-linear phenomena here, we instead refer the reader to the recent work of Zemskova and Grisouard (2022). Here, the energetics of a steady rotating flow over tall topography is analysed in detail, and used to help explain observations in the Southern Ocean by Cusack et al. (2020).
@@ 4.4 Steady flow in three dimensions
Figure 7: Topographic stress for three-dimensional steady flows with high Froude number $(Fr\gg 1)$. In panels (a)–(c), the drag from the simulation data is compared with both a linear $Nh_{0}^{2}WU$ and quadratic $h_{0}WU^{2}$ scaling. The default parameters are as in Table 1. Note that in panels (a)–(d) the uncertainty in the computed stress was always less than 10% so the small error bars have been omitted.
For steady flow tests in three-dimensions, we again compare simulations with low Froude numbers $Fr\lesssim 1$ to the Bell (1975) theory. However, for tests with high Froude numbers $Fr\gg 1$, it does not make sense to compare our results directly to KLP2010 as this theory was only derived in the two-dimensional case. Thus, for high Froude numbers, we instead focus on how the stress scales with different parameters.
Figure 6 shows a comparison between the stress from our simulations (denoted $F_{\text{3d}}$) and $F_{\text{Bell3d}}$ when the parameters $h_{0}$, $N$, $U$, $f$ and $W$ are varied. For tests varying the hill height $h_{0}$ and buoyancy frequency $N$ (Figure 6a) we find, similar to the two-dimensional case, that $F_{\text{Bell3d}}$ underestimates the stress but still gives a mostly accurate scaling (i.e. $\text{Stress}\propto Nh_{0}^{2}$). However, for tests that vary the flow speed $U$, Coriolis parameter $f$ and hill width $W$ (Figure 6b), we see that $F_{\text{Bell3d}}$ significantly underestimates the stress as the Rossby number
$Ro\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{U}{|f|W},$ (36)
@@
where $\tilde{K}=W\sqrt{k^{2}+\ell^{2}}$ and $(kW,\ell W)=(\tilde{K}\cos\theta,\tilde{K}\sin\theta)$. From (4.4) we see that as $Ro$ increases so too does the lee wave stress. Therefore, a possible explanation for the discrepancy between $F_{\text{Bell3d}}$ and the simulation data is that as $Ro$ decreases, fewer lee waves are formed and the stress is instead generated by other non-linear motions such as partial blocking (see e.g., Trossman et al., 2016; Klymak et al., 2021). To parametrise this phenomenon, we note that from Figure 6b the ratio $F_{\text{Bell3d}}/F_{3d}$ has a clear linear dependence on the ratio $|f|W/U$. In particular, our data points give a correlation coefficient between $F_{\text{Bell3d}}/F_{3d}$ and $|f|W/U$ of $R^{2}=0.89$. This suggests that we should use the scaling
@@
Figure 7 shows the results of our tests for the high Froude number $Fr\gg 1$ regime. In this case, we find that the stress deviates significantly from the approximate $Nh_{0}^{2}WU$ scaling found in other regimes. Instead we find that the stress obeys a quadratic law, being proportional to $h_{0}WU^{2}$, similar to what one expects in the atmosphere (e.g., Lott and Miller, 1997). In fact, the simple parametrisation, which sets the stress equal to $h_{0}WU^{2}$ (with proportionality constant $1$) works remarkably well for our choice of parameters. However, as is usual with quadratic drag laws, it is likely that a different proportionality constant would be required if the shape of the topography differed substantially from a Gaussian hill. The stress then appears to be mostly independent of $N$ and $f$ (Figures 7d and 7e); only increasing or decreasing by small amounts as these parameters are varied. We note however, that for larger values of $f$, our computed value of the stress has a large uncertainty. These results can be directly compared to recent work of Klymak et al. (2021) which involved a similar analysis but with many, randomly distributed obstacles. Klymak et al. also observed this $f$ independence for the stress but an approximately linear dependence on $N$ for high Froude numbers. Therefore, we do not expect the stress to remain independent of $N$ for more complicated non-isolated topography.
@@ 4.5 Mixed flow in two dimensions
We now analyse our results for mixed flows with a horizontal velocity of the form (20) and stress of the form (31). In what follows we let $\widetilde{F}_{m}$ and $\widetilde{F}_{p}$ be the values for $F_{m}$ and $F_{p}$ if $U_{\text{tidal}}$ or $U_{\text{mean}}$ are set to be zero respectively. In Figure 8 we show our computed values of $F_{m}/\widetilde{F}_{m}$ and $F_{p}/\widetilde{F}_{p}$ for a range of parameters. This lets us discern the impact of a tidal flow on a steady flow and vice versa. When the Froude number is small ($Fr\lesssim 1$) these ratios are also compared to the expected ratio from Bell’s parametrisations defined in Sections 22.2 and 22.3 (Figures 8e and 8f).
@@
For low Froude numbers (Figures 8e and 8f) we also observe a dependence of $F_{p}/\widetilde{F}_{p}$ on the mean flow speed $U_{\text{mean}}$ and the hill width $W$. In particular, $F_{p}/\widetilde{F}_{p}$ slightly exceeds 1 as $U_{\text{mean}}$ increases (Figure 8e), and significantly exceeds $1$ for small $W$ (Figure 8f). This is likely linked to the fact that the classical lee wave frequency $kU_{\text{mean}}$ depends explicitly on $U_{\text{mean}}$ and implicitly on $W$. In particular, with the Fourier convention in (6), the wavenumber spectrum for a Gaussian hill $h(x)=h_{0}e^{-x^{2}/(2W^{2})}$ is
@@
In terms of developing a refined parametrisation, one could fit a linear trend line for $F_{p}/\widetilde{F}_{p}$ against $U_{\text{mean}}$. However, we are not sure how useful this would be since the increase in $F_{p}/\widetilde{F}_{p}$ is only small with $U_{\text{mean}}$ and it is clear that $W$ is the far more sensitive parameter. It appears difficult however, to develop a more accurate parametrisation for small $W$, especially given how non-linear the trend is between $F_{p}/\widetilde{F}_{p}$ and $W$ in Figure 8f. Notably, low values of $W$ correspond to large values of the steepness parameter (see (29)), so that non-linear effects are expected to take place in this regime.
@@ 4.6 Mixed flow in three dimensions
Finally we consider mixed flows in three dimensions, with our results for this regime shown in Figure 9. Note that we use the same notation as in Section 44.5). Like the two-dimensional case, we also find that $F_{m}/\widetilde{F}_{m}\approx 1$ and $F_{p}/\widetilde{F}_{p}\approx 1$ for the majority of tests. In fact, this approximation appears to hold true even more than in the two-dimensional case (cf. Figure 8).
Of particular note is the far less dramatic increase in $F_{p}/\widetilde{F}_{p}$ near the critical latitude $|f|\approx\omega$. This behaviour seems primarily related to our previous observations for purely tidal flows, rather than some new phenomena specifically linked to three-dimensional mixed flows. Namely, in two-dimensions $\widetilde{F}_{p}$ essentially reaches 0 at the critical latitude (Figure 2c) causing the periodic component of the mixed (and thus Doppler-shifted) stress $F_{p}$ to be very large by comparison. Then, by contrast, in three dimensions the value of $\widetilde{F}_{p}$ does not approach zero when $|f|=\omega$ (Figure 4c), resulting in a more subdued peak for the ratio $F_{p}/\widetilde{F}_{p}$.
Another notable observation in Figure 9 is the distinct increase in the ratio $F_{m}/\widetilde{F}_{m}$ as $h_{0}$ increases (Figure 9c), which was not observed in the two-dimensional setting. To try to understand this increase, we recall the results from our purely steady flow tests with large Froude number in three dimensions (see Figure 7). For this regime, the mean stress followed a quadratic drag law, being proportional to $h_{0}WU^{2}$ as opposed to the $h_{0}^{2}WU$ scaling for low Froude numbers. Thus, averaging over one tidal period $T$, one expects that for large Froude numbers
$\frac{F_{m}}{\widetilde{F_{m}}}$ $=\frac{1}{h_{0}WU_{\text{mean}}^{2}T}\int_{0}^{T}h_{0}W\left[U_{\text{mean}}+U_{\text{tidal}}\cos(\omega t)\right]^{2}\mathrm{d}t$ (41)
$=1+\frac{U_{\text{tidal}}^{2}}{2U_{\text{mean}}^{2}}.$ (42)
For the case $U_{\text{tidal}}=U_{\text{mean}}=0.1$ represented in Figure 9c, this gives $F_{m}/\widetilde{F}_{m}=1.5$, which is not too far off the model output for the larger values of $h_{0}$. In future work, it would be useful to understand this phenomenon further. In particular, it would be worthwhile evaluating the formula (42) throughout a wider parameter space. Moreover, it would also be useful to better understand the location and behavior of the transition between the low and high Froude number regime. Based on the limited data in Figure 4 this transition appears around $Fr\approx 10-20$, and corresponds to larger uncertainty in $F_{m}/\widetilde{F}_{m}$.
@@
$F_{3d}=F_{m}+F_{p}\cos(\omega t+\phi),$ (43)
with $\phi\approx 0$. By comparison, in two dimensions we found that $\phi$ was much larger as $|f|$ increased (Figure 8g). Such an effect is also in contrast to the purely tidal flow case, whereby the phase $\phi$ approaches $\pi/2$ for large $|f|$.
@@ 5 Discussion and summary of suggested parametrisations
In the previous section, we tested existing topographic stress parametrisations for oceanic flows against a suite of simulations. Particular parametrisations of interest were those of Jayne and St. Laurent (2001) and Shakespeare et al. (2020) for tidal flows, as well as Bell (1975) and Klymak et al. (2010) for steady flows. Spanning a wide parameter space for an isolated Gaussian hill, we found that in many cases these parametrisations proved accurate, in many cases even when tested beyond the validity of their underlying assumptions. However, there were also regimes in which these existing parametrisations fell short, suggesting the need for an alternate or modified parametrisation.
We also studied the case of a mixed flow in Sections 44.5 and 44.6. Here, we primarily compared our simulation data against the naive parametrisation which assumes that the tidal and steady components of a flow are independent. Our data yielded the cases for which this naive parametrisation was accurate, and cases for which a more sophisticated parametrisation needs to be developed.
@@ 5.1 Suggested tidal flow parametrisations
For a purely tidal flow, we found that the parametrisations $F_{\text{SAH2d}}$ (Equation (2.1.2)) and $F_{\text{SAH3d}}$ (Equation (2.1.2) due to Shakespeare et al. (2020, 2021a) were very accurate, mainly just requiring a small modification for large hill heights. In particular, for a purely oscillatory (tidal) flow in two dimensions, we suggest setting
@@
noting that this parametrisation may overestimate the stress when $h_{0}\rightarrow H$.
@@
for large hill widths $W$, where $F_{\text{SAH3d}}$ is as in (2.1.2). From our testing, W=10000mW=10000\ $\mathrm${m} was a sufficiently large width to notice this effect (with $h_{0}=100\ \mathrm{m}$). For more modest values of $W$, we instead suggest setting $F_{3d}=F_{\text{SAH3d}}$.